OpenStudy (anonymous):

PLEASE HELP!!!!!!! I NEED HELP ON THIS QUESTION!!!! Darnell is selecting names from a hat to determine who will choose the family activities for the weekend. The names in the hat include his mother, father, sister, and him. Calculate the probability of each compound event. TERMS: A)P(Darnell and sister); without replacement. B)P(Mother and Father); with replacement. C)P(Parent and sister); with replacement. D)P(Parent and child); without replacement. ASWER TYPES: A) one sixth B) one twelfth C) one third D) one eighth

OpenStudy (btaylor):

A) 1/4 x 1/3 = 1/12 B) 1/4 x 1/4 = 1/16 C) 1/2 x 1/4 = 1/8 D) 1/2 x 1/2 = 1/4

OpenStudy (anonymous):

THANKS!!!!!!!!!!!!!